ROTATIONAL SPECTRA OF CYCLOPROPYLMETHYL GERMANE AND CYCLOPROPYLMETHYL SILANE: DIPOLE MOMENT AND BARRIER TO METHYL GROUP ROTATION

\underline{REBECCA A. PEEBLES}, SEAN A. PEEBLES, MICHAEL D. FOELLMER, JONATHAN M. MUR-RAY, MICHAL M. SERAFIN and AMANDA L. STEBER, Department of Chemistry, Eastern Illinois University, 600 Lincoln Avenue, Charleston, IL 61920; GAMIL A. GUIRGIS and RICHARD LIB-ERATORE, Department of Chemistry and Biochemistry, The College of Charleston, Charleston, SC 29424; JAMES R. DURIG and CHARLES J. WURREY, Department of Chemistry, University of Missouri - Kansas City, Kansas City, MO 64110.

The  Fourier-transform  microwave  spectra  correspond-
ing  to  the  gauche  conformer  of  five  isotopologues

of cyclopropylmethyl germane and three isotopologues of cyclopropylmethyl silane have been assigned. a-, b- and c-type transitions are all split into doublets by internal rotation of the methyl group; a global fit of the A and E state spectra leads to barriers to rotation of 4.736(6) kJ mol^{−1} in the germane and ca. 6.7 kJ mol^{−1} in the silane. Stark effect measurements reveal similar dipole moment components for both molecules: μ_{a} = 0.1782(10) D, μ_{b} = 0.581(4) D, μ_{c} = 0.305(9) D and μ_{total} = 0.680(5) D for the germane and preliminary values of μ_{a} = 0.196(4) D, μ_{b} = 0.670(5) D, μ_{c} = 0.363(8) D and μ_{total} = 0.787(6) D for the silane.

Nuclear quadrupole coupling constants for the ^{73}Ge isotopologue (I = \frac{9}{2} ) of cyclopropylmethyl germane have been determined and are in good agreement with values predicted at the B3LYP/6-311++G(3df,3pd) level. The electric field gradient at the Ge nucleus will be compared with related germanes and the barriers to rotation for both the germane and the silane will be compared with similar species.